# An Exploration to the Correlation Structure and Clustering of Macroeconomic Variables 11footnotemark: 122footnotemark: 2

### v2

@@ 3 Clustering on MEVs
    We select a list of target MEVs, denoted as $X$. They are generally those MEVs believed to be relevant to our portfolio. For this project we select 44 US national level monthly MEVs as listed in Table 2 in A.
@@
    Given the fact that most MEVs are levels and not stationary, we focus on their stationary transformations so the mean and correlation are meaningful and internally consistent. We denote them as $f(X)$. For this project, we use 3 months, 6 months and 12 months simple differences for all rates like MEVs and $\log$ differences for the rest. So there are eventually $33\times 4$ = 132 MEVs transformations.
@@
    For the above $f(X)$, e.g., 132 MEV transformations here, for any interested window of time $T_{i}$, we apply the MEVs clustering on the realization of $f(X)$ during $T_{i}$, with the consistent stopping criteria. We denote this clustering operation as $\mathcal{C}(f(X),T_{i})$. In this project, $T_{i}$s are the annual two-year rolling windows from 2003 to 2022, i.e., $T_{1}$ is 2003-01 to 2004-12, $T_{2}$ is 2004-01 to 2005-12, …, $T_{19}$ is 2021-01 to 2022-12. Thus, $f(X)$ is a 24 (months) by 132 (transformations) matrix for each $T_{i}$. For every two-year rolling window, we first standardize the 132 transformations and then apply clustering. We use a simple stopping criteria $\lambda_{2}<1$ as mentioned in Section 2.3.3. Note that since we have 12 months log or simple differences transformation in $f(X)$, each $T_{i}$ actually uses three years economic data. We use this setup to sync with the 2-3 years crisis window in both 2008 GFC and various hypothetical stress scenarios.
@@
From this visualization it is apparent that we tend to get more clusters in benign years and less clusters in adverse years. The minimum number of clusters, 15, is achieved during the 2007-2008 GFC and the 2020-2021 COVID-19 period. For curiosity, we applied the same clustering on four different hypothetical scenarios coming with CCAR2023 exercise, in the 2-year window from 2023-01 to 2024-12999In the CCAR2023 exercise, the launch point, i.e., the last actual observation date, is 2022-12-31. MEV curves after this date are hypothetical., with the results presented in Table 1.
@@ 4 Summary and Conclusion
Appendix A 44 MEVs List
Here is the 44 MEVs used in Section 3.
Category MEVs
General Economy US Real GDP
US Nominal GDP
US real PCE levels
US nominal PCE levels
Real DIsposable Income
Nominal Disposable Income
Nominal Personal Income
National Median Income
US CPI Inflation Index level
Real Durable Goods Spending
Real Non-Durable Goods Spending
Retail Sales
Housing Market Housing Starts
Housing Price Index (HPI)
Commerical Real Estate Price Index
US Population
Labor Market US Unemployment Rate
Nonfarm Employment
Rates USD Libor 1M/3M/6M/1Y
US Treasury Yield 3M/6M/2Y/3Y/5Y/7Y/10Y/20Y/30Y
USD Interest Rate Swap Bond Yield 2/3/4/5/6/7/10/15/20/30Y
BBB Coporate Yield
Spread AA Coporate Spread
BBB Coporate Bond Spred
Table 2: 44 MEVs List

### v3

@@ 3 Clustering on MEVs
    We select a list of target MEVs, denoted as $X$. They are generally those MEVs believed to be relevant to our portfolio.
    For this project we select 44 US national level monthly MEVs from five categories, i.e., general economy, housing market, labor market, rates and spread, including the 16 key MEVs provided by FRB, i.e., Real and Nominal GDP, Real and Nominal Income, unemployment, CPI, 3M/5Y/10Y Treasury Yield, BBB yield, Mortgage rate, Dow Jones Total Stock Market Index, HPI, CRE Price Index and VIX.
@@
    Given the fact that most MEVs are levels and not stationary, we focus on their stationary transformations so the mean and correlation are meaningful and internally consistent. We denote them as $f(X)$.
    For this project, we use 3 months, 6 months and 12 months simple differences for all rates like MEVs and $\log$ differences for the rest. So there are eventually $44\times 3$ = 132 MEVs transformations.
@@
    For the above $f(X)$, e.g., 132 MEV transformations here, for any interested window of time $T_{i}$, we apply the MEVs clustering on the realization of $f(X)$ during $T_{i}$, with the consistent stopping criteria. We denote this clustering operation as $\mathcal{C}(f(X),T_{i})$.
    In this project, $T_{i}$s are the annual two-year rolling windows from 2003 to 2022, i.e., $T_{1}$ is 2003-01 to 2004-12, $T_{2}$ is 2004-01 to 2005-12, …, $T_{19}$ is 2021-01 to 2022-12. Thus, $f(X)$ is a 24 (months) by 132 (transformations) matrix for each $T_{i}$.
    For every two-year rolling window, we first standardize the 132 transformations and then apply clustering. We use a simple stopping criteria $\lambda_{2}<1$ as mentioned in Section 2.3.3.
    Note that since we have 12 months log or simple differences transformation in $f(X)$, each $T_{i}$ actually uses three years economic data. We use this setup to sync with the 2-3 years crisis window in both 2008 GFC and various hypothetical stress scenarios.
@@
From this visualization it is apparent that we tend to get more clusters in benign years and less clusters in adverse years. The minimum number of clusters, 15, is achieved during the 2007-2008 GFC and the 2020-2021 COVID-19 period. For curiosity, we applied the same clustering on two different hypothetical scenarios coming with CCAR2023 exercise, in the 2-year window from 2023-01 to 2024-12999In the CCAR2023 exercise, the launch point, i.e., the last actual observation date, is 2022-12-31. MEV curves after this date are hypothetical., with the results presented in Table 2.
@@
We further tried CCAR2022 and get relatively similar results: For both Baseline and Severly Adverse scenario, the number of clusters is 2 less than it is in CCAR2023.
Hypothetical Scenarios Number of Clusters
FRB Baseline 19
FRB Severely Adverse 13
Table 2: CCAR2022 Hypothetical Scenarios 2-Year Clustering Results
Our speculation to this observation, i.e., in CCAR2022 both the Baseline and Severely Adverse scenario have 2 less clusters than they have in CCAR2023, is because when FRB was constructing the CCAR2022 scenarios in 2021, the baseline and outlook was rather pessimistic, possibly because of the rampaging covid; while in 2022 everything was stabilized so FRB was less pessimistic when constructing the CCAR2023 scenarios and thus they have 2 more clusters.